Peter is forty years old and his son is ten years old. How long will it take so that will Peter be twice as old as his son?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52800)

.

Let x be the number of years to pass.


Then your basic equation is this

    40 + x = 2*(10+x)


Simplify and find x


    40 + x = 20 + 2x

    40 - 20 = 2x - x

      20    =    x


ANSWER.  In 20 years.


CHECK.  In 20 years their ages will be  10+20 = 30 (the son)  and  40 + 20 = 60  (the father).

Solved.



Answer by Alan3354(69443)

Peter is forty years old and his son is ten years old. How long will it take so that will Peter be twice as old as his son?
--------------
The difference remains 30 years.
---> ages 30 and 60